OpenStudy (peterg2001):

Which of the following sets of ordered pairs is not a direct variation? (0, 0); (-2, 4); (3, -6) (1, 1); (-2, -2); (3, 3) (10, 2); (15, 3); (20, 4) (0, 0); (1, 3); (2, 4) @lizz123 @chrisk123 @DanJS

OpenStudy (anonymous):

1 i think but im not sure you might want to ask around

OpenStudy (danjs):

y = k*x use one point from each set (x,y) and figure the k value out for that set, then test the other two points to make sure they work

OpenStudy (peterg2001):

A B C Or D

OpenStudy (danjs):

for example the second one. y=k*x and point (1,1) 1 = k*1 k = 1/1 = 1 y = 1*x test (-2,-2) and (3,3) to make sure they work in y=1*x

OpenStudy (danjs):

you kiddin me ABC or D?

OpenStudy (peterg2001):

hey i realy need answer nOW